Please can you help me? Amanda has 500 in ten dollar, five dollar, and one dollar bills. There were 100 bills in all, and 20 more five dollar bills than ten dollars bills. How many one-dollar bills did Amanda have? Round to nearest whole number
Answer by scott8148(6628)

let x=tens, so x+20=fives and 100-(2x+20)=ones

10x+5(x+20)+1(100-(2x+20))=500